# Moving through Cartesian products, coronas and joins in general position

### v1

@@ 4.1 Corona product graphs
We first recall that the general position number of corona product graphs was studied in [14]. We next bound the value of the mobile general position number of the corona product $G\odot H$.
@@
When both $G$ and $H$ are complete graphs, $G\odot H$ is a block graph, hence the following consequence can also be deduced from [16, Theorem 2.3].
@@
If $r,s\geq 1$, then, $\operatorname{Mob_{gp}}(K_{r}\odot K_{s})=\max\{r,s+1\}$.
Note that Corollary 4.2 demonstrates the sharpness of all the bounds in Theorem 4.1. For another sharpness example, in which the upper and lower bounds do not coincide, consider $G=K_{2}$ and $H=C_{4}$. Then we have that $\operatorname{Mob_{gp}}(K_{2})=2$, $\operatorname{Mob_{gp}}(C_{4}\vee K_{1})=2$, and $\operatorname{gp}(C_{4}\vee K_{1})=3$. It can be noted that $\operatorname{Mob_{gp}}(K_{2}\odot C_{4})=3=\operatorname{gp}(C_{4}\vee K_{1})$. For another infinite family, let $G$ be an arbitrary tree, and $H$ the edgeless graph of order at least two. Since in this case $G\odot H$ is a tree, we have $\operatorname{Mob_{gp}}(G\odot H)=2$ (see [16, Theorem 2.3]), which is also the value of the lower bound in Theorem 4.1.
@@
Starting from some fixed $Y_{i}$, we move the robot $R$ placed at $\left(\left\lceil\frac{n}{2}\right\rceil+i\right)^{\prime}$ through the vertices of the set $[\left(\left\lceil\frac{n}{2}\right\rceil+i\right),(i-1)]$ and their attached leaves. For $20\leq j\leq\left\lfloor\frac{n}{2}\right\rfloor-$ 2, let $Z_{j}$ be the sequence of three moves
$\left(\left\lceil\frac{n}{2}\right\rceil+j\right)^{\prime}\rightsquigarrow\left(\left\lceil\frac{n}{2}\right\rceil+j\right)\rightsquigarrow\left(\left\lceil\frac{n}{2}\right\rceil+j+1\right)\rightsquigarrow\left(\left\lceil\frac{n}{2}\right\rceil+j+1\right)^{\prime}.$
Perform the series of moves $Z_{0},Z_{1},\dots,Z_{\left\lfloor\frac{n}{2}\right\rfloor-2}$ in that order. Observe that after these moves we are left with the robots occupying the set $Y_{i-1}$ and at each stage the robots remained in general position, since the paths between robots in $Y_{i}\setminus\{\left(\left\lceil\frac{n}{2}\right\rceil+i\right)^{\prime}\}$ pass through $[i,\left\lceil\frac{n}{2}\right\rceil+i-1]$. Therefore starting at $Y_{0}$ and repeating this process for $i=0,-1,\dots,-\left\lceil\frac{n}{2}\right\rceil$ results in all vertices being visited by a robot and $Y_{0}$ is a mobile general position set as claimed.
We now prove that $\operatorname{Mob_{gp}}(C_{n}\odot K_{1})\leq\left\lceil\frac{n}{2}\right\rceil+1$. Note that each set $\{i,i^{\prime}\}$, $i\in\mathbb{Z}_{n}$, is a maximal general position set, so we must have $|X\cap\{i,i^{\prime}\}|\leq 1$ for each $i\in\mathbb{Z}_{n}$ and any $\operatorname{Mob_{gp}}$-set $X$. Consider the moment that a robot visits the vertex 00. Let $j,k$ be the smallest and largest values in $[1,n-1]$, respectively, such that there is a robot in $\{j,j^{\prime}\}$ and $\{k,k^{\prime}\}$. To avoid the robot in $\{k,k^{\prime}\}$ having a shortest path to the robot in $\{j,j^{\prime}\}$ through the robot at 00 we must have $k-j\leq\left\lceil\frac{n}{2}\right\rceil-1$. As each set $\{i,i^{\prime}\}$ contains at most one robot for $i\in[j,k]\cup\{0\}$ and no robots for $i\in[k+1,n-1]\cup[1,j-1]$, this gives an upper bound of $\left\lceil\frac{n}{2}\right\rceil+1$ robots in the graph. ∎
@@ 4.2 Joins of graphs
For the lower bound, assume that $\omega(H)\leq\omega(G)$. We can start with robots at a maximum clique $W_{H}$ of $H$ and one robot in $G$. The robot in $G$ can visit every vertex of $G$, since during this process the vertices occupied form a clique in $G\vee H$. At the end, this robot moves into a maximum clique $W_{G}$ of $G$. After that, all the robots from $W_{H}$ but one move into $W_{G}$. At that time, only one robot remains in $H$ and, by the same argument, it can visit every vertex of $H$.
@@
Note that if $G$ and $H$ are both triangle-free graphs, then the upper bound and the lower bound of Theorem 4.4 coincide when neither graph is $K_{1}$. For a triangle-free graph $G$, $\operatorname{Mob_{gp}}(G\vee K_{1})$ could be either two or three. It is easily seen that for cycles we have $\operatorname{Mob_{gp}}(C_{4}\vee K_{1})=2$ and $\operatorname{Mob_{gp}}(C_{n}\vee K_{1})=3$ for $n\geq 5$. The first example shows that the lower bound for $\operatorname{Mob_{gp}}(G\vee K_{1})$ is tight, but it is an open question whether this can happen for graphs with large clique number.
@@
If $G$ and $H$ are triangle-free and both have order at least two, then $\operatorname{Mob_{gp}}(G\vee H)=3$.
To show that the upper bound is tight, consider the join $K_{r}^{-}\vee K_{s}^{-}$, where $K_{n}^{-}$ represents a complete graph minus one edge. If $x_{1},x_{2}$ is the pair of non-adjacent vertices in $K_{r}^{-}$ and $y_{1},y_{2}$ is the pair of non-adjacent vertices of $K_{s}^{-}$, then $(V(K_{r}^{-})\setminus\{x_{2}\})\cup(V(K_{s}^{-})\setminus\{y_{1},y_{2}\})$ is a mobile general position set, as the set of occupied vertices forms a clique in $K_{r}^{-}\vee K_{s}^{-}$ and the robot at $x_{1}$ can follow the route $x_{1}\rightsquigarrow y_{1}\rightsquigarrow x_{2}\rightsquigarrow y_{2}$ to visit the remaining vertices. This matches the upper bound. More generally, the same argument works when $G$ and $H$ are both joins of cliques with empty graphs.
To demonstrate sharpness of the lower bound, for $r\geq 2$ take the join $K_{r}\vee K^{+}_{r+1}$, where $K^{+}_{r+1}$ is the complete graph $K_{r+1}$ with an added leaf $x$. Suppose that $r+2$ robots can traverse this graph in general position, and focus on the moment that there is a robot at $x$. As $r+2$ robots cannot be stationed on $K_{r+1}^{+}$, there must be a robot on $K_{r}$ and the positions occupied on $K_{r}$ and $K_{r+1}^{+}$ must both induce cliques. Hence every vertex of $K_{r}$ must contain a robot and in $K_{r+1}^{+}$ there is a robot at $x$ and its support vertex $x^{\prime}$. However, there are no legal moves in this configuration. Thus, $\operatorname{Mob_{gp}}(K_{r}\vee K^{+}_{r+1})\leq r+1=\min\{\omega(K_{r}),\omega(K^{+}_{r+1})\}+1$.
For an example of graphs with arbitrarily large clique number that meet the upper bound in $\operatorname{Mob_{gp}}(G\vee K_{1})\leq\omega(G)+1$, consider the birdcage graph $B_{n}$ formed as follows. Let $U$ be a clique on vertices $\{u_{1},\dots,u_{n}\}$, $V$ be an empty graph on vertices $\{v_{1},\dots,v_{n}\}$ and $z$ an additional vertex, and add edges $u_{i}\sim v_{i}$ and $v_{i}\sim z$ for $i\in[n]$. The graph $B_{n}$ has clique number $n$ and we show that $n+1$ robots can traverse $B_{n}\vee K_{1}$ in general position. Denote the vertex of the $K_{1}$ by $x$. Start with robots at $U\cup\{x\}$ and make the move $x\rightsquigarrow z$. Then for each $i\in[n]$ make the two moves $u_{i}\rightsquigarrow v_{i}\rightsquigarrow u_{i}$. It is easily seen that the robots are always in general position and visit all the vertices of $B_{n}\vee K_{1}$.
@@ 5 Concluding remarks and open problems
  • We have seen that $\operatorname{Mob_{gp}}(P_{\infty}\operatorname{\,\square\,}P_{\infty})=4=\operatorname{gp}(P_{\infty}\operatorname{\,\square\,}P_{\infty})$. In [18, Theorem 1] it was proven that $\operatorname{gp}(P_{\infty}^{k,\Box})=2^{2^{k-1}}$, where $P_{\infty}^{k,\Box}$ the $k$-tuple Cartesian of the infinite path $P_{\infty}$. It is therefore of interest to determine whether $\operatorname{Mob_{gp}}(P_{\infty}^{k,\Box})=2^{2^{k-1}}$ holds for larger values of $k$.
@@ Acknowledgments
Sandi Klavžar was supported by the Slovenian Research and Innovation Agency (ARIS) under the grants P1-0297, N1-0355, and N1-0285. D. Kuziak and I. G.Yero have been partially supported by the Spanish Ministry of Science and Innovation through the grant PID2023-146643NB-I00. D. Kuziak acknowledges the support from “Ministerio de Educación, Cultura y Deporte”, Spain, under the “José Castillejo” program for young researchers (reference number: CAS22/00081) to make a temporary visit to the University of Ljubljana, where this investigation has been partially developed. The research of James Tuite was partially supported by a grant from the Crowther Fund.

### v2

@@ 4.1 Corona product graphs
The first paper [17] on the mobile general position problem briefly considered mobile general position sets in rooted products. This suggests investigating the problem in corona products, which can also be viewed as a kind of rooted product. The general position number of corona product graphs was studied in [14]. We now bound the value of the mobile general position number of the corona product $G\odot H$.
@@
When both $G$ and $H$ are complete graphs, $G\odot H$ is a block graph, hence the following consequence can also be deduced from [17, Theorem 2.3].
@@
If $r,s\geq 1$, then $\operatorname{Mob_{gp}}(K_{r}\odot K_{s})=\max\{r,s+1\}$.
Note that Corollary 4.2 demonstrates the sharpness of all the bounds in Theorem 4.1. For another sharpness example, in which the upper and lower bounds do not coincide, consider $G=K_{2}$ and $H=C_{4}$. Then we have that $\operatorname{Mob_{gp}}(K_{2})=2$, $\operatorname{Mob_{gp}}(C_{4}\vee K_{1})=2$, and $\operatorname{gp}(C_{4}\vee K_{1})=3$. It can be noted that $\operatorname{Mob_{gp}}(K_{2}\odot C_{4})=3=\operatorname{gp}(C_{4}\vee K_{1})$. For another infinite family, let $G$ be an arbitrary tree, and $H$ the edgeless graph of order at least two. Since in this case $G\odot H$ is a tree, we have $\operatorname{Mob_{gp}}(G\odot H)=2$ (see [17, Theorem 2.3]), which is also the value of the lower bound in Theorem 4.1.
@@
Starting from some fixed $Y_{i}$, we move the robot $R$ placed at $\left(\left\lceil\frac{n}{2}\right\rceil+i\right)^{\prime}$ through the vertices of the set $\left(\left\lceil\frac{n}{2}\right\rceil+i\right),(i-1)]$ and their attached leaves, and finally leave the robot at $(i-1)^{\prime}$. After these moves we are left with the robots occupying the set $Y_{i-1}$ and at each stage the robots remained in general position, since the paths between robots in $Y_{i}\setminus\{\left(\left\lceil\frac{n}{2}\right\rceil+i\right)^{\prime}\}$ pass through $i,\left\lceil\frac{n}{2}\right\rceil+i-1]$. Therefore starting at $Y_{0}$ and repeating this process for $i=0,-1,\dots,-\left\lceil\frac{n}{2}\right\rceil$ results in all vertices being visited by a robot and $Y_{0}$ is a mobile general position set as claimed.
We now prove that $\operatorname{Mob_{gp}}(C_{n}\odot K_{1})\leq\left\lceil\frac{n}{2}\right\rceil+1$. Note that each set $\{i,i^{\prime}\}$, $i\in\mathbb{Z}_{n}$, is a maximal general position set, so we must have $|X\cap\{i,i^{\prime}\}|\leq 1$ for each $i\in\mathbb{Z}_{n}$ and any $\operatorname{Mob_{gp}}$-set $X$. Consider the moment that a robot visits the vertex 0. Let $j,k$ be the smallest and largest values in [1,n1][1,n-1], respectively, such that there is a robot in $\{j,j^{\prime}\}$ and $\{k,k^{\prime}\}$. To avoid the robot in $\{k,k^{\prime}\}$ having a shortest path to the robot in $\{j,j^{\prime}\}$ through the robot at 0 we must have $k-j\leq\left\lceil\frac{n}{2}\right\rceil-1$. As each set $\{i,i^{\prime}\}$ contains at most one robot for $i\in[j,k]\cup\{0\}$ and no robots for $i\in[k+1,n-1]\cup[1,j-1]$, this gives an upper bound of $\left\lceil\frac{n}{2}\right\rceil+1$ robots in the graph. ∎
@@ 4.2 Joins of graphs
For the lower bound, assume that $\omega(H)\leq\omega(G)$. We can start with robots at a maximum clique $W_{H}$ of $H$ and one robot in $G$. The robot in $G$ can visit every vertex of $G$, since during this process the occupied vertices always form a clique in $G\vee H$. At the end, this robot moves into a maximum clique $W_{G}$ of $G$. After that, all the robots from $W_{H}$ but one move into $W_{G}$. At that time, only one robot remains in $H$ and, by the same argument, it can visit every vertex of $H$.
@@
Note that if $G$ and $H$ both have clique number two, then the upper and lower bounds of Theorem 4.4 coincide. For a triangle-free graph $G$, $\operatorname{Mob_{gp}}(G\vee K_{1})$ could be either two or three. It is easily seen that for cycles we have $\operatorname{Mob_{gp}}(C_{4}\vee K_{1})=2$ and $\operatorname{Mob_{gp}}(C_{n}\vee K_{1})=3$ for $n\geq 5$. The first example shows that the lower bound for $\operatorname{Mob_{gp}}(G\vee K_{1})$ is tight, but it is an open question whether this can happen for graphs with large clique number.
@@
If $G$ and $H$ both have clique number two, then $\operatorname{Mob_{gp}}(G\vee H)=3$.
To show that the upper bound in Theorem 4.4 is tight, consider the join $K_{r}^{-}\vee K_{s}^{-}$, where $K_{n}^{-}$ represents a complete graph minus one edge. If $x_{1},x_{2}$ is the pair of non-adjacent vertices in $K_{r}^{-}$ and $y_{1},y_{2}$ is the pair of non-adjacent vertices of $K_{s}^{-}$, then $(V(K_{r}^{-})\setminus\{x_{2}\})\cup(V(K_{s}^{-})\setminus\{y_{1},y_{2}\})$ is a mobile general position set, as the set of occupied vertices forms a clique in $K_{r}^{-}\vee K_{s}^{-}$ and the robot at $x_{1}$ can follow the route $x_{1}\rightsquigarrow y_{1}\rightsquigarrow x_{2}\rightsquigarrow y_{2}$ to visit the remaining vertices. This matches the upper bound. More generally, the same argument works when $G$ and $H$ are both joins of cliques with empty graphs.
To demonstrate sharpness of the lower bound, for $r\geq 2$ take the join $K_{r}\vee K^{+}_{r+1}$, where $K^{+}_{r+1}$ is the complete graph $K_{r+1}$ with an added leaf $x$. Suppose that a set of at least $r+2$ robots can traverse this graph in general position, and focus on the moment that there is a robot at $x$. As $r+2$ robots cannot be stationed on $K_{r+1}^{+}$, there must be a robot on $K_{r}$ and the positions occupied on $K_{r}$ and $K_{r+1}^{+}$ must both induce cliques. Hence every vertex of $K_{r}$ must contain a robot and in $K_{r+1}^{+}$ there is a robot at $x$ and its support vertex $x^{\prime}$. However, there are no legal moves in this configuration. Thus, $\operatorname{Mob_{gp}}(K_{r}\vee K^{+}_{r+1})\leq r+1=\min\{\omega(K_{r}),\omega(K^{+}_{r+1})\}+1$.
For an example of graphs with arbitrarily large clique number that meet the upper bound in $\operatorname{Mob_{gp}}(G\vee K_{1})\leq\omega(G)+1$, consider the birdcage graph $B_{n}$ formed as follows. Let $U$ be a clique on vertices $\{u_{1},\dots,u_{n}\}$, $V$ be an empty graph on vertices $\{v_{1},\dots,v_{n}\}$ and an additional vertex $z$, and add edges $u_{i}\sim v_{i}$ and $v_{i}\sim z$ for $i\in[n]$. The graph $B_{n}$ has clique number $n$ and we show that $n+1$ robots can traverse $B_{n}\vee K_{1}$ in general position. Denote the vertex of the $K_{1}$ by $x$. Start with robots at $U\cup\{x\}$ and make the move $x\rightsquigarrow z$. Then for each $i\in[n]$ make the two moves $u_{i}\rightsquigarrow v_{i}\rightsquigarrow u_{i}$. It is easily seen that the robots are always in general position and visit all the vertices of $B_{n}\vee K_{1}$.
@@ 5 Concluding remarks and open problems
  • We have seen that $\operatorname{Mob_{gp}}(P_{\infty}\operatorname{\,\square\,}P_{\infty})=4=\operatorname{gp}(P_{\infty}\operatorname{\,\square\,}P_{\infty})$. In [19, Theorem 1] it was proven that $\operatorname{gp}(P_{\infty}^{k,\Box})=2^{2^{k-1}}$, where $P_{\infty}^{k,\Box}$ is the $k$-tuple Cartesian product of the infinite path $P_{\infty}$. It is therefore of interest to determine whether $\operatorname{Mob_{gp}}(P_{\infty}^{k,\Box})=2^{2^{k-1}}$ holds for larger values of $k$.
@@ Acknowledgments
Sandi Klavžar was supported by the Slovenian Research and Innovation Agency (ARIS) under the grants P1-0297, N1-0355, and N1-0285. D. Kuziak and I. G. Yero have been partially supported by the Spanish Ministry of Science and Innovation through the grant PID2023-146643NB-I00. D. Kuziak acknowledges the support from “Ministerio de Educación, Cultura y Deporte”, Spain, under the “José Castillejo” program for young researchers (reference number: CAS22/00081) to make a temporary visit to the University of Ljubljana, where this investigation has been partially developed. The research of James Tuite was partly carried out during a visit to University of Ljubljana supported by a grant from the Crowther Fund. Ethan Shallcross conducted this research with funding from an Open University research bursary. The authors thank the anonymous referees for their suggestions.
Conflicts of interest
Sandi Klavžar is an Associate Editor of the Discrete Applied Mathematics journal and was not involved in the review and decision-making process of this article. The authors do not have any additional conflicts of interest to declare.